# Gradient-based Sample Selection for Faster Bayesian Optimization

### v2

@@ 4 Bayesian Optimization with Gradient-based Sample Selection
4.1 Gradient Information with GP
In the previous section, we introduced how to use the gradient to select a diverse subset of samples. However, within the Bayesian optimization framework using Gaussian processes, we are more concerned with the impact on the GP posterior of the information carried by the observation $y_{i}$. Therefore, in light of the pilot estimate-based gradient information acquisition method in [44], we propose a new method for gradient information acquisition in GPs: $g_{i}\;=\;\frac{\partial}{\partial y_{i}}\,\log p(\mathbf{y}\mid\mathbf{X},\theta).$ We focus on the gradient of output $\mathbf{y}$ rather than input $\mathbf{X}$ as the derivative of $\mathbf{y}$ is simpler and directly reflects each sample’s contribution. In a GP model, given a set of samples $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ that follows a multivariate normal distribution with mean $\mu$ and covariance matrix $\mathbf{K}$, where $\mathbf{K}$ is constructed from a kernel function $k(\mathbf{x}_{i},\mathbf{x}_{j};\theta)$ and ${\theta}$ represents the hyperparameters, the probability density function of a multivariate Gaussian distribution is $p(\mathbf{y}|\mathbf{X},{\theta})=\frac{1}{(2\pi)^{n/2}|\mathbf{K}|^{1/2}}\exp\left(-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\right)$. Taking the logarithm of it, we derive the log-likelihood function:
$\begin{split}\log p(\mathbf{y}|\mathbf{X},{\theta})=-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log(2\pi).\end{split}$ (2)
@@ 4.1 Gradient Information with GP
The log-likelihood function in (2) comprises three terms. The first term, $-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}$, represents the sample fit under the covariance structure specified by $\mathbf{K}$. The second term, $-\frac{1}{2}\log|\mathbf{K}|$, penalizes model complexity through the log-determinant of the covariance matrix. The third term, $-\frac{n}{2}\log(2\pi)$, is a constant to the parameters and thus does not affect the gradient calculation.
The derivative of log likelihood with respect to $\mathbf{y}$ directly measures how sensitive this log-likelihood is to each observation $y_{i}$. Intuitively, if changing $y_{i}$ significantly alters the value of (2), that sample has a large marginal contribution to the fit. Hence, in subset selection schemes, one can use these gradient magnitudes to gauge how important each sample is, potentially adjusting their weights or deciding which samples to retain in a subset. To define the gradient for each sample, we first quantify each sample’s contribution to the log-likelihood. Since the second and third terms in (2), $-\frac{1}{2}\log|\mathbf{K}|$ and $-\frac{n}{2}\log(2\pi)$, do not depend on $\mathbf{y}$, both have no contribution to the gradient. Consequently, the gradient of the log-likelihood with respect to $\mathbf{y}$ is given by: $\frac{\partial\log p(\mathbf{y|X},{\theta})}{\partial\mathbf{y}}=-\mathbf{K}^{-1}(\mathbf{y}-\mu)$. Note that the $i$-th component, $-(\mathbf{K}^{-1}(\mathbf{y}-\mu))_{i}$, corresponds to the partial derivative of the log-likelihood with respect to $\mathbf{y}_{i}$. Thus, we define the gradient $g_{i}$ for each sample $\mathbf{y}_{i}$ as:
$g_{i}=\frac{\partial\log p(\mathbf{y|X},{\theta})}{\partial\mathbf{y}}=-(\mathbf{K}^{-1}(\mathbf{y}-\mu))_{i}.$ (3)
After obtaining the gradient information of each sample, we can use the maximization of the variance of the angle between the two gradients as an alternative goal (i.e. maintaining the diversity of gradient directions). This gradient calculation is computationally efficient as the value of $\mathbf{K}^{-1}$ in (3) is available while updating the GP. Furthermore, the complexity of the additional computational burden introduced by the gradient calculation is $O(n^{2})$, and it is negligible compared to the complexity of GP updates (which is $O(n^{3})$), especially when $n$ is large.
@@ 4.3 Gradient-based Sample Selection BO
Detailed Implementations. In the initialization phase, $n_{0}$ initial samples $\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$ are observed; and the initial sample set $D$, the buffer size $M$, and total budget $N$ are specified. In each iteration, the GP posterior is updated on the current sample set $D$, an acquisition function (e.g., UCB) is built to select the next point $x_{t}$, the corresponding observation $y_{t}=f(x_{t})$ is obtained, and the sample $(x_{t},y_{t})$ is added to $D$. To manage each iteration’s computational complexity, a buffer check and gradient-based selection step are performed. Specifically, if the current size of $D$ is less than or equal to $M$, the GP is updated using all samples. Otherwise, a gradient-based sample selection step is performed to identify a set of the most representative samples. Note that the $n_{0}$ initial samples and the newly acquired sample $(x_{t},y_{t})$ are always added into the subset, as they provide base information for the GP model and ensure that recently observed information is always retained, respectively. Besides the initial $n_{0}$ and the newly observed samples, $(M-n_{0}-1)$ samples are selected by minimizing the sum of pairwise cosine similarities among their gradients. The resulting subset $\mathcal{U}$, containing $M$ samples, is then used to fit the GP model. The complete procedure is outlined in Algorithm 4.3.
@@
(1) Dynamic Buffer Size. In practice, the buffer size should be prespecified by the users. However, the value is often unavailable in advance. Instead, we propose a dynamic adjustment mechanism to determine the buffer size. We define a tolerable maximum factor $Z$ to accelerate GP computations. Let $\bar{T}$ be the average time for a single initial iteration and $T_{\text{current}}$ be the current iteration’s computation time. If $T_{\text{current}}$ exceeds the user-specified threshold $Z\times\bar{T}$, the buffer size is set to be the number of all current samples, i.e., $M=|D|$. This adaptive strategy ensures that the algorithm balances computational efficiency with the goal of utilizing as much data as possible, thereby maintaining high predictive accuracy without incurring excessive costs.
@@
(2) Preserving Initial Samples and Latest Observations. During the procedure, the initial $n_{0}$ samples are always included in the subset. These samples capture essential information about the overall function landscape, and good initialization samples give us useful information about the global situation. Additionally, the newly acquired sample, $(x_{t},y_{t})$, is also included in the subset. This ensures that the GP model incorporates the latest data, maintaining its relevance and accuracy. Consequently, the algorithm prevents valuable information from being prematurely excluded. Additionally, this essentially alleviates a limitation of sparse GP in BO [17]: the constrained representation size may hinder the full integration of new observations into the model.
(3) Random Perturbation to Escape Local Optima. Subset selection may cause being trapped in the local optima, as it iteratively selects a subset of “locally optimal” samples. This issue arises when certain samples are frequently chosen by the acquisition function but consistently excluded from the sample subset, hindering the exploration of other valuable samples. To address this issue, Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,(\sigma_{d})^{2})$ is added to each computed gradient $g_{i}$, resulting in perturbed gradients $\tilde{g}_{i}=g_{i}+\epsilon_{i}$. This perturbation introduces variability into the selection process, helping the algorithm escape local optima and facilitating the exploration of a more diverse set of samples.
@@
1:  Initialization: Obtain $n_{0}$ initial samples $D=\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$, and fit an initial GP model. Set buffer size $M>n_{0}$, total budget $N$, average initial iteration time $\bar{T}$, and threshold factor $Z$. Initialize switched $\leftarrow$ false.
2:  for $t=n_{0}+1$ to $N$ do
@@
13:        Form subset $\mathcal{U}$ by forcing in the $n_{0}$ initial samples and the newest sample $(x_{t},y_{t})$.
14:        Select the remaining $M-n_{0}-1$ samples from $D$ via the gradient-based criterion.
@@ 5 Theoretical Analysis
Gaussian Process Upper Confidence Bound (GP-UCB [46]) is a popular algorithm for sequential decision-making problems. We propose an extension to GP-UCB by incorporating gradient-based sampling. Some assumptions and notations follow the previous proof. Here we analyze the error of the subset fitted GP and prove that the regret of the GSSBO with GP-UCB algorithm is bounded.
@@
$|\Delta\mu(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,|\Delta\sigma^{2}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,$ (4)
@@
Assumption 5.1. Assume there exist constants $a$, $b$, and $L$ such that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies a Lipschitz continuity condition, providing confidence bounds on the derivatives of the GP sample paths $f$: $P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,d.$ A typical example of such a kernel is the squared exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2l^{2}}\right)$, where $l$ is the length-scale parameter and $\sigma^{2}$ represents the noise variance. Then we propagate the error in Theorem 5.1 through the GP posterior to bound the GSSBO regret.
@@
[…]
@@
The sketch proof for the main theorem is relegated to the Appendix. The main theoretical challenge lies in evaluating the error between the low-rank approximation $\hat{\mathbf{K}}$ and the full $\mathbf{K}_{\mathcal{D}\mathcal{D}}$. By invoking spectral norm inequalities and the Nyström approximation theory, $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ can be bounded. We prove that select $\mathcal{U}$ by maximizing gradient-direction diversity coincides with the greedy Nyström column-selection, then derive select $\mathcal{U}$ by maximizing gradient-direction diversity is better than random selection in standard Nyström approximation. Then, we merge the resulting linear and $\beta$-scaled error terms into a single penalty in the UCB construction. We also determine the smallest subset size $M_{\min}$. This establishes a regret bound for GSSBO, which is similar to that of classical GP-UCB. From a practical relevance perspective, Theorem 5.1 indicates that limiting the GP to a smaller, well-chosen subset does not substantially degrade posterior accuracy in either the mean or variance estimates. Restricting the subset size $M$ confers significant computational savings while ensuring performance closely matches that of a standard GP-UCB using all samples. Of note, we also observe that, compared with the classical UCB results, our GSSBO retains the same fundamental structure of an upper confidence bound approach. Still, it restricts the GP fitting to a gradient-based sample subset, lowering computational costs.
@@ 6 Experiments
Experimental Setup.
We choose UCB as the acquisition function in GSSBO, and compare GSSBO with the following benchmarks: (1) Standard GP-UCB [46]; (2) Random Sample Selection GP-UCB (RSSBO), which mirrors our approach in restricting the sample set size but chooses which samples randomly; (3) VecchiaBO [12], which utilizes the Vecchia approximation method in BO; (4) LR-First m [14], which uses a low-rank approximation based on the first $m$ samples. (5) SparseGP [15], which introduces a small number of inducing points to obtain a sparse approximation of GP; (6) IPA [24], which multiplies an expected-improvement-based quality function by the GP posterior and selects inducing points regarding quality-diversity. and (7) mini-META [25] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold. (More baseline can be found in the appendix.) We employ a Matérn $5/2$ kernel for the GP, with hyperparameters learned via maximum likelihood estimation. Both GSSBO and RSSBO use the same buffer size $M$, dynamically adjusted by a parameter $Z=4$, $M$ is usually around 100 and will change with $Z$. The gradient-perturbation noise is set to $\sigma^{2}=0.01$. The size of the initial set is 20. Each experiment is repeated 50 times, and the total number of iterations is 1000. All experiments were conducted on a MacBook Pro with Apple M2 Pro (10-core CPU, 16 GB unified RAM).
6.1 Synthetic Test Problems
@@
Optimization Performance Analysis. Figure 2 compares methods on multiple functions, evaluating cumulative regret. Overall, GSSBO achieves comparable performance with the Standard GP-UCB while outperforming the RSSBO. In low-dimensional problems such as Eggholder2 and Hart6, GSSBO has a subtle gap with Standard GP-UCB and VecchiaBO. In contrast, GSSBO significantly outperforms other methods. In high-dimensional settings, such as on Levy20, Powell50, and Rastrigin100, GSSBO achieves the smallest cumulative regret, surpassing other methods. From the experimental results, we can observe that the cumulative regret of our algorithm is sublinear, which is consistent with the theoretical results. In particular, GSSBO achieves these results with a significant reduction in computation time, as shown in Figure 3. GSSBO strikes a combination of performance and efficiency in scalable optimization tasks by maintaining near-baseline regret while significantly improving computational efficiency.

### v3

@@ 4 Bayesian Optimization with Gradient-based Sample Selection
4.1 Gradient Information from GP
In the previous section, we introduced how to use the gradient to select a diverse subset of samples. However, within the Bayesian optimization framework using Gaussian processes, we do not have access to second-order information with respect to the input $x_{i}$, nor an explicit differentiable form of the objective to build $x_{i}$-gradients or sensitivities. We are more concerned with the impact on the GP posterior of the information carried by the observation $y_{i}$. Therefore, in light of the pilot estimate-based gradient information acquisition method in [Zhu, 2016], we propose a new method for gradient information acquisition in GPs: $g_{i}\;=\;\frac{\partial}{\partial y_{i}}\,\log p(\mathbf{y}\mid\mathbf{X},\theta).$ We focus on the gradient of output $\mathbf{y}$ rather than input $\mathbf{X}$ as the derivative of $\mathbf{y}$ is simpler and directly reflects each sample’s contribution. In a GP model, given a set of samples $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ that follows a multivariate normal distribution with mean $\mu$ and covariance matrix $\mathbf{K}$, where $\mathbf{K}$ is constructed from a kernel function $k(\mathbf{x}_{i},\mathbf{x}_{j};\theta)$ and $\theta$} represents the hyperparameters, the probability density function of a multivariate Gaussian distribution is $p(\mathbf{y}|\mathbf{X},{\theta})=\frac{1}{(2\pi)^{n/2}|\mathbf{K}|^{1/2}}\exp\left(-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\right)$. Taking the logarithm of it, we derive the log-likelihood function:
$\begin{split}\log p(\mathbf{y}|\mathbf{X},{\theta})=&-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\\ &-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log(2\pi).\end{split}$ (2)
@@ 4.1 Gradient Information from GP
The log-likelihood function in (2) comprises three terms. The first term, $\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}$, represents the sample fit under the covariance structure specified by $\mathbf{K}$. The second term, $\frac{1}{2}\log|\mathbf{K}|$, penalizes model complexity through the log-determinant of the covariance matrix. The third term, $\frac{n}{2}\log(2\pi)$, is a constant to the parameters and thus does not affect the gradient calculation.
The derivative of log likelihood with respect to $\mathbf{y}$ directly measures how sensitive this log-likelihood is to each observation $y_{i}$. Intuitively, if changing $y_{i}$ significantly alters the value of (2), that sample has a large marginal contribution to the fit. Hence, in subset selection schemes, one can use these gradient magnitudes to gauge how important each sample is, potentially adjusting their weights or deciding which samples to retain in a subset. To define the gradient for each sample, we first quantify each sample’s contribution to the log-likelihood. Since the second and third terms in (2), $\frac{1}{2}\log|\mathbf{K}|$ and $\frac{n}{2}\log(2\pi)$, do not depend on $\mathbf{y}$, both have no contribution to the gradient. Consequently, the gradient of the log-likelihood with respect to $\mathbf{y}$ is given by: $\frac{\partial\log p(\mathbf{y|X},{\theta})}{\partial\mathbf{y}}=-\mathbf{K}^{-1}(\mathbf{y}-\mu)$. Note that the $i$-th component, $\big(\mathbf{K}^{-1}(\mathbf{y}-\mu)\big)_{i}$, corresponds to the partial derivative of the log-likelihood with respect to $\mathbf{y}_{i}$; we denote this scalar sensitivity by $g_{i}$. Thus, we further define the vector embedding for each sample $i$,
$\mathbf{g}_{i}\;\triangleq\;\frac{\partial}{\partial y_{i}}\!\left(\frac{\partial\log p(\mathbf{y|X},\theta)}{\partial\mathbf{y}}\right)=-\,\mathbf{K}^{-1}\mathbf{e}_{i},$ (3)
where $\mathbf{e}_{i}$ is the $i$-th standard basis vector. After obtaining the gradient information of each sample, we can use the maximization of the variance of the angle between the two gradients as an alternative goal (i.e. maintaining the diversity of gradient directions). This gradient calculation is computationally efficient as the value of $\mathbf{K}^{-1}$ is available while updating the GP. Furthermore, the complexity of the additional computational burden introduced by the gradient calculation is $O(n^{2})$, and it is negligible compared to the complexity of GP updates (which is $O(n^{3})$), especially when $n$ is large.
@@ 4.3 Gradient-based Sample Selection BO
Detailed Implementations. In the initialization phase, $n_{0}$ initial samples $\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$ are observed; and the initial sample set $D$, the buffer size $M$, and total budget $N$ are specified. In each iteration, the GP posterior is updated on the current sample set $D$, an acquisition function (e.g., UCB) is built to select the next point $x_{t}$, the corresponding observation $y_{t}=f(x_{t})$ is obtained, and the sample $(x_{t},y_{t})$ is added to $D$. To manage each iteration’s computational complexity, a buffer check and gradient-based selection step are performed. Specifically, if the current size of $D$ is less than or equal to $M$, the GP is updated using all samples. Otherwise, a gradient-based sample selection step is performed to identify a set of the most representative samples. Note that the newly acquired sample $(x_{t},y_{t})$ are always added into the subset, as they provide base information for the GP model and ensure that recently observed information is always retained, respectively. Besides the newly observed samples, (M1)(M-1) samples are selected by minimizing the sum of pairwise cosine similarities among their gradients. The resulting subset $\mathcal{U}$, containing $M$ samples, is then used to fit the GP model. The complete procedure is outlined in Algorithm 4.3.
@@
(1) Dynamic Buffer Size. In practice, the buffer size should be prespecified by the users. However, the value is often unavailable in advance. Instead, we propose a dynamic adjustment mechanism to determine the buffer size. We define a tolerable maximum factor $Z$ to accelerate GP computations. Let $\bar{T}$ be the average wall-clock time for a single initial iteration and $T_{\text{current}}$ be the current iteration’s computation time. If $T_{\text{current}}$ exceeds the user-specified threshold $Z\times\bar{T}$, the buffer size is set to be the number of all current samples, i.e., $M=|D|$. This adaptive strategy ensures that the algorithm balances computational efficiency with the goal of utilizing as much data as possible, thereby maintaining high predictive accuracy without incurring excessive costs.
@@
(2) Preserving Latest Observations. During the procedure, the newly acquired sample, $(x_{t},y_{t})$, is also included in the subset. This ensures that the GP model incorporates the latest data, maintaining its relevance and accuracy. Consequently, the algorithm prevents valuable information from being prematurely excluded. Additionally, this essentially alleviates a limitation of sparse GP in BO [McIntire et al., 2016]: the constrained representation size may hinder the full integration of new observations into the model. This will also help to escape local optima from iteratively selecting a subset of “locally optimal” samples, since this observed sample will not be observed in next iteration.
@@
1:Initialization: Obtain $n_{0}$ initial samples $D=\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$, and fit an initial GP model. Set buffer size $M>n_{0}$, total budget NN, average initial iteration time $\bar{T}$, and threshold factor $Z$. Initialize switched $\leftarrow$ false.
2:for $t=n_{0}+1$ to NN do
@@
13:   Form subset $\mathcal{U}$ by forcing in $(x_{t},y_{t})$.
14:   Select the remaining M1M-1 samples from $D$ via the gradient-based criterion.
@@ 5 Theoretical Analysis
Gaussian Process Upper Confidence Bound (GP-UCB [Srinivas et al., 2009]) is a popular algorithm for sequential decision-making problems. We propose an extension to GP-UCB by incorporating gradient-based sampling. Some assumptions and notations follow  [Srinivas et al., 2009]. Here we analyze the error of the subset fitted GP and prove that the regret of the GSSBO with GP-UCB algorithm is bounded.
@@
[…]
@@
Assumption 1. Assume there exist constants $a$, $b$, and $L$ such that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies a Lipschitz continuity condition, providing confidence bounds on the derivatives of the GP sample paths $f$: $P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,d.$ A typical example of such a kernel is the squared exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2l^{2}}\right)$, where $l$ is the length-scale parameter and $\sigma^{2}$ represents the noise variance. Then we propagate the error in Theorem 5.1 through the GP posterior to bound the GSSBO regret.
@@
(Regret Bound for GSSBO with UCB) Let $\mathcal{X}\subseteq[0,r]^{d}$ be compact and convex, $d\in\mathbb{N},r>0$, let $A=\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|\;C_{M}\,(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2})$, $B_{n}=\|\mathbf{k}_{*\mathcal{D}}\|\;\sqrt{\,C_{M}\;\Bigl(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr)}$, where $\lambda_{M+1}$ is the (M+1)(M+1)‐th largest eigenvalue of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$, $\epsilon_{g}$ is the Nyström approximation error parameter for gradient based sample selection, and $\sigma_{\min}>0$ satisfies $\sigma(x)\geq\sigma_{\min}$ for all $x\in\mathcal{X}$, where $\sigma(x)$ is posterior standard deviation. Under Assumption 1, for any arbitrarily small $\delta\in(0,1)$, choose $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl[2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right)\Bigr]\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ Where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}\left(\sqrt{dN\gamma_{T_{N}}}\right)$. Where $\gamma_{T_{N}}$ is the information gain. Specifically, with $C_{1}=\frac{8}{\log(1+\sigma^{-2})},$ we have:
@@
The sketch proof for the main theorem is relegated to the Appendix. The main theoretical challenge lies in evaluating the error between the low-rank approximation $\hat{\mathbf{K}}$ and the full $\mathbf{K}_{\mathcal{D}\mathcal{D}}$. By invoking spectral norm inequalities and the Nyström approximation theory, $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ can be bounded. We prove that select $\mathcal{U}$ by maximizing gradient‐direction diversity coincides with the greedy Nyström column‐selection, then derive select $\mathcal{U}$ by maximizing gradient‐direction diversity is better than random selection in standard Nyström approximation. Then, we merge the resulting linear and $\beta$-scaled error terms into a single penalty in the UCB construction. We also determine the smallest subset size $M_{\min}$. This establishes a regret bound for GSSBO, which is similar to that of classical GP-UCB. From a practical relevance perspective, Theorem 5.1 indicates that limiting the GP to a smaller, well-chosen subset does not substantially degrade posterior accuracy in either the mean or variance estimates. Restricting the subset size $M$ confers significant computational savings while ensuring performance closely matches that of a standard GP-UCB using all samples. Of note, we also observe that, compared with the classical UCB results, our GSSBO retains the same fundamental structure of an upper confidence bound approach. Still, it restricts the GP fitting to a gradient-based sample subset, lowering computational costs.
@@ 6 Experiments
Experimental Setup.
We choose UCB as the acquisition function in GSSBO, and compare GSSBO with the following benchmarks: (1) Standard GP-UCB [Srinivas et al., 2009]; (2) Random Sample Selection GP-UCB (RSSBO), which mirrors our approach in restricting the sample set size but chooses which samples randomly; (3) VecchiaBO [Jimenez and Katzfuss, 2023], which utilizes the Vecchia approximation method in BO; (4) LR-First m [Williams and Seeger, 2000], which uses a low-rank approximation based on the first $m$ samples. (5) SparseGP [Lawrence et al., 2002], which introduces a small number of inducing points to obtain a sparse approximation of GP; (6) IPA [Moss et al., 2023], which multiplies an expected‐improvement‐based quality function by the GP posterior and selects inducing points regarding quality‐diversity. and (7) mini-META [Calandriello et al., 2022] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold. (More baseline and high dimensional task can be found in the appendix.) We employ a Matérn $5/2$ kernel for the GP, with hyperparameters learned via maximum likelihood estimation. Both GSSBO and RSSBO use the same buffer size $M$, dynamically adjusted by a parameter $Z=4$, $M$ is usually around 100 and will change with $Z$. The gradient-perturbation noise is set to $\sigma^{2}=0.01$. The size of the initial set is 20. Each experiment is repeated 50 times, and the total number of iterations is 1000. All experiments were conducted on a MacBook Pro with Apple M2 Pro (10-core CPU, 16 GB unified RAM).
6.1 Synthetic Test Problems
Optimization Performance Analysis. Figure 2 compares methods on multiple functions, evaluating cumulative regret. Overall, GSSBO achieves comparable performance with the Standard GP-UCB while outperforming the RSSBO. In low-dimensional problems such as Eggholder2 and Hart6, GSSBO has a subtle gap with Standard GP-UCB and VecchiaBO. In contrast, GSSBO significantly outperforms other methods. In high-dimensional settings, such as on Levy20, Powell50, and Rastrigin100, GSSBO achieves the smallest cumulative regret, surpassing other methods. From the experimental results, we can observe that the cumulative regret of our algorithm is sublinear, which is consistent with the theoretical results. In particular, GSSBO achieves these results with a significant reduction in computation time, as shown in Figure 3. GSSBO strikes a combination of performance and efficiency in scalable optimization tasks by maintaining near-baseline regret while significantly improving computational efficiency. The reason GP‑UCB outperforms most other baselines in Figure 2 is that it employs full-data standard GP updates, whereas the other baselines use approximate algorithms that incur information loss, leading to their inferior performance.